C(x) = e^0.1x + 500 , where x is the number of items produced, what is the total cost of producing 70 units?
Answer by RAY100(1637)

c(x)=e^.1x+500
for x=70
c(70)=e^7+500
c(70) =1096.63+500=1596.63